# Marginally interpretable spatial logistic regression with bridge processes

### v1

@@ Abstract
In including random effects to account for dependent observations, the odds ratio interpretation of logistic regression coefficients is changed from population-averaged to subject-specific. This is unappealing in many applications, motivating a rich literature on methods that maintain the marginal logistic regression structure without random effects, such as generalized estimating equations. However, for spatial data, random effect approaches are appealing in providing a full probabilistic characterization of the data that can be used for prediction. We propose a new class of spatial logistic regression models that maintain both population-averaged and subject-specific interpretations through a novel class of bridge processes for spatial random effects. These processes are shown to have appealing computational and theoretical properties, including a scale mixture of normal representation. The new methodology is illustrated with simulations and an analysis of childhood malaria prevalence data in the Gambia.
@@ 1 Introduction
We propose a new class of marginally interpretable spatial logistic regression models based on a novel spatial bridge process. We identify the normal-scale mixture representation of the bridge distribution and propose its multivariate extension. In contrast to existing copula-based constructions that have bridge-distributed marginals (Parzen et al.,, 2011; Boehm et al.,, 2013; Swihart et al.,, 2014), the bridge process has appealing properties, such as transparent correlation structure and comes with significant computational benefits.
@@ 3.1 Partially collapsed Gibbs sampler
Under a Bayesian framework, we describe a posterior inference procedure for the proposed model (4) under an independent prior $p(\beta,\phi)=p(\beta)p(\phi)$ based on Markov chain Monte Carlo. For simplicity, we focus on a normal prior for $p(\beta)$ and a known correlation kernel $\mathcal{R}$, but the extension to inference under a normal scale mixture shrinkage prior or inference on unknown parameter(s) of $\mathcal{R}$ are straightforward by adding additional sampling steps.
@@ 3.2 Scalable computation with low-rank dependence structure
The Algorithm 1 involves several inversions and determinant calculations of the $n\times n$ matrices, which creates a computational bottleneck when the number of spatial locations $n$ is large. In such cases, we can adopt a correlation kernel that has a low-rank structure. For example, ~$\tilde{\mathcal{R}}(s,s^{\prime})=r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})+1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\}$ (Finley et al.,, 2009) where $\{\tilde{s}_{1},\dots,\tilde{s}_{q}\}$ are knots with $q\ll n$, $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}\in\mathbb{R}^{q}$ and $R_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a $q\times q$ matrix. Using the Woodbury matrix identity and the matrix determinant lemma, the computation related to the inversion and calculation of the determinant of $n\times n$ matrices can be reduced to those of $q\times q$ matrices; see Appendix A.2 for details.
4 Simulation study
@@
We perform simulation studies to compare the proposed approach with existing methods for either marginal or conditional inference. As a marginal competitor, we consider the spatial generalized estimating equation method based on pairwise log-odds ratios (Cattelan and Varin,, 2018). As a conditional competitor, we consider a Bayesian logistic mixed model with Gaussian process random effects. For both bridge and Gaussian process random effect models, we choose Cauchy priors for the regression coefficients and half-Cauchy priors for the random effect standard deviation, following Gelman, (2006); Gelman et al., (2008).
@@
We generate data based on the spatial logistic model (1) under two different processes with bridge-distributed marginals, one with the proposed model and another based on the Gaussian copula process with bridge marginals. This ensures both population-averaged effects $\beta^{M}=\phi\beta$ and site-specific effects $\beta$ are well-defined, where we set $\phi=0.7$. We choose $250$ spatial locations uniformly at random from unit square domain $\mathscr{S}=[0,1]^{2}$, among which $50$ locations are held out for assessing predictive performance. We use Matérn correlation kernel $\mathcal{R}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. The same correlation kernel is applied to the latent process in the Gaussian copula. We set $N_{i}=10$ for all locations and set $p=2$ including intercept, where the non-intercept covariate $\{x_{ij}\}$ is generated from an independent standard normal distribution. The true fixed-effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$. This data generation process is repeated 200 times.
@@
Table 1: Comparison of population-averaged $(\hat{\beta}_{1}^{M})$ and site-specific $(\hat{\beta}_{1})$ estimates across three methods along with coverage probabilities at nominal level 0.95.
Data from bridge process Data from Gaussian copula process
Population-average Site-specific Population-average Site-specific
$\rho$ Method Sp GEE BrP RE BrP RE GP RE Sp GEE BrP RE BrP RE GP RE
0.05 Bias 0.015 0.005 0.004 0.002 0.008 0.006 0.001 -0.001
RMSE 0.176 0.128 0.068 0.068 0.063 0.051 0.063 0.063
Coverage 40.0% 96.5% 96.5% 95.5% 94.5% 100.0% 98.5% 98.5%
0.1 Bias 0.033 -0.009 0.001 -0.001 0.026 -0.021 0.001 -0.002
RMSE 0.184 0.140 0.074 0.074 0.084 0.069 0.065 0.065
Coverage 48.7% 95.0% 94.5% 93.5% 92.0% 100.0% 98.5% 97.0%
@@
Sp GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP RE, proposed bridge process random effect model; GP RE, Gaussian process random effect model; RMSE, root mean squared error.
@@
The simulation results are summarized in Table 1 in terms of population-averaged estimate $\hat{\beta}^{M}_{1}$ and site-specific estimate $\hat{\beta}_{1}$, where posterior mean estimates are used for Bayesian approaches. All methods yield consistent estimates in general, but exhibit some differences in terms of mean square error and coverage probabilities depending on underlying data-generating models.
@@
Under the bridge process data-generating model, the spatial generalized estimating equation method gives higher mean squared error and narrower confidence intervals of $\hat{\beta}^{M}_{1}$, leading to low coverage probability. Here the robust standard error correction cannot be easily applied due to the lack of independent replicates; see Cattelan and Varin, (2018) for a discussion. The proposed model performs well for both $\hat{\beta}^{M}_{1}$ and $\hat{\beta}_{1}$ as the model is correctly specified, and the Gaussian process random effect model gives a similar result to the proposed model.
@@
Under the Gaussian copula process data-generating model, the spatial generalized estimating equation method provides reasonable coverage probabilities compared to the proposed model, yet has a slightly higher mean square error. The proposed model gives a wider credible interval for $\hat{\beta}^{M}_{1}$, and a similar pattern can be observed for $\hat{\beta}_{1}$ but to a milder degree. In all settings, the mean square error of estimators increases as spatial dependence becomes stronger. Detailed simulation settings and further comparison summaries of $\hat{\beta}^{M}_{0},\hat{\beta}_{0}$ and predictive performance are available in Appendix A.4.
5 Childhood malaria prevalence in the Gambia

### v2

@@ Abstract
In including random effects to account for dependent observations, the odds ratio interpretation of logistic regression coefficients is changed from population-averaged to subject-specific. This is unappealing in many applications, motivating a rich literature on methods that maintain the marginal logistic regression structure without random effects, such as generalized estimating equations. However, for spatial data, random effect approaches are appealing in providing a full probabilistic characterization of the data that can be used for prediction. We propose a new class of spatial logistic regression models that maintain both population-averaged and subject-specific interpretations through a novel class of bridge processes for spatial random effects. These processes are shown to have appealing computational and theoretical properties, including a scale mixture of normal representation. The new methodology is illustrated with simulations and an analysis of childhood malaria prevalence data in Gambia.
@@ 1 Introduction
We propose a new class of marginally interpretable spatial logistic regression models based on a novel spatial bridge process. We identify the normal-scale mixture representation of the bridge distribution and propose its multivariate extension. In contrast to existing copula-based constructions that have bridge-distributed marginals (Lin et al.,, 2010; Li et al.,, 2011; Parzen et al.,, 2011; Boehm et al.,, 2013; Swihart et al.,, 2014), the bridge process has appealing properties, such as transparent correlation structure and comes with significant computational benefits.
@@ 3.1 Partially collapsed Gibbs sampler
Many existing inference methods for spatial logistic regression models, or in general for spatial generalized linear mixed models, are based on Gaussian process spatial random effects. For example, integrated nested Laplace approximation (Rue et al.,, 2009; Lindgren et al.,, 2011) handles a non-Gaussian response variable with Laplace approximation together with latent Gaussian structure to carry out fast approximate inference. When random effects no longer follow Gaussian processes to preserve population-averaged interpretation of coefficients, existing approaches cannot be easily applied; instead, based on Markov chain Monte Carlo, generic sampling algorithms such as Metropolis-Hastings are often adopted (Gory et al.,, 2021). However, such an algorithm is sensitive to tuning and is not appropriate for sampling high-dimensional spatial random effects.
Under a Bayesian framework, we describe how the posterior inference for the proposed model (4) can be efficiently carried out, leveraging normal mixture representation of the bridge distribution. For simplicity, we focus on prior $p(\beta,\phi)=p(\beta)p(\phi)$ with a normal prior for $\beta$, some proper prior on $\phi\in(0,1)$, and a known correlation kernel $\mathcal{R}$, but the extension to inference under a normal scale mixture shrinkage prior or inference on unknown parameter(s) of $\mathcal{R}$ are straightforward by adding additional sampling steps.
@@ 3.2 Scalable computation with low-rank dependence structure
The Algorithm 1 involves several inversions and determinant calculations of the $n\times n$ matrices, creating a computational bottleneck when the number of spatial locations $n$ is large. Specifically, step 1 involves the inversion of a $n\times n$ matrix and step 2 involves the evaluation of the density and sampling of the multivariate normal distribution of dimension $n$. Without special structures in the corresponding covariance or precision matrices, the computation becomes prohibitive as the number of spatial locations $n$ increases.
Due to the normal mixture representation, several existing computational strategies for Gaussian processes can be easily integrated into bridge processes. One way is to introduce a low-rank structure on the correlation kernel, following the strategy of Finley et al., (2009) for Gaussian processes. Given a positive definite correlation kernel $\mathcal{R}$, denote $R_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ as a $q\times q$ matrix formed from $q$ knot locations $\tilde{s}_{k}\in\mathscr{S}$, $k=1,\dots,q$. Then, we consider a correlation kernel $\tilde{\mathcal{R}}$ defined as
~$\tilde{\mathcal{R}}(s,s^{\prime})=r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})+1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\},$ (8)
where $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}$ is a $q\times 1$ vector and the term $1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\}$ in (8) ensures that ~$\tilde{\mathcal{R}}(s,s)=1$. Using the Woodbury matrix identity and the matrix determinant lemma, the computation related to the inversion and calculation of the determinant of $n\times n$ matrices can be reduced to those of $q\times q$ matrices, which gives huge computational benefits when $q\ll n$; see Appendix A.2 for details. In light of Proposition 2, the seamless incorporation of scalable Gaussian process methods through a hierarchical formulation is a distinctive feature of bridge processes.
4 Simulation studies
4.1 Comparison with existing approaches
@@
First, we perform a simulation study to compare the proposed approach with existing methods for either marginal or conditional inference. As a marginal competitor, we consider the spatial generalized estimating equation method based on pairwise log-odds ratios (Cattelan and Varin,, 2018). As a conditional competitor, we consider a Bayesian logistic mixed model with Gaussian process random effects.
@@
We generate data based on the spatial logistic model (1) under two different processes with bridge-distributed marginals, one with the proposed model and another based on the Gaussian copula process with bridge marginals. This ensures both population-averaged effects $\beta^{M}=\phi\beta$ and site-specific effects $\beta$ are well-defined, where we set $\phi=0.7$. We choose spatial locations uniformly at random from unit square domain $\mathscr{S}=[0,1]^{2}$ to decide training and test location with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})=(200,50)$, where $50$ test locations are held out for assessing predictive performance. We use Matérn correlation kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. The same correlation kernel is applied to the latent process in the Gaussian copula. We set $N_{i}=10$ for all locations and set $p=2$ including the intercept, where the non-intercept covariate $\{x_{ij}\}$ is generated from an independent standard normal distribution. The true fixed-effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$. This data generation process is repeated 200 times.
We describe details of prior specifications. Following Gelman et al., (2008), we choose Cauchy priors for $\beta$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 1.25 for $\beta_{1}$. For the random effect standard deviation $\sigma_{u}$, we choose a weakly informative half-Cauchy prior (Gelman,, 2006), namely $p(\sigma_{u})=2/\{\pi(1+\sigma_{u}^{2})\}$. For the bridge process models where $\sigma_{u}=3^{-1/2}\pi(\phi^{-2}-1)^{1/2}$, this becomes $p(\phi)=(12)^{1/2}/[\{\pi^{2}-(\pi^{2}-3)\phi^{2}\}(1-\phi^{2})^{1/2}]$ for $\phi\in(0,1)$. For the prior of spatial range $\rho$ in Matérn correlation kernel with known smoothness 1.5, we assign uniform prior $\rho\sim\mathrm{Unif}(0.001,0.3)$ for both bridge and Gaussian process random effect models, which corresponds to effective range approximately between 0.00475 and 1.425 in unit square domain.
Next, we describe inference settings. For the spatial generalized estimating equation method of Cattelan and Varin, (2018) with the same notations therein, we used the number of bins $B=13$, radius $h=0.05$, and $d_{\mathrm{max}}=0.3$ to obtain the empirical spatial lorelogram, and optimize the parameters $\alpha_{2}$ (sill) and $\alpha_{3}$ (range) of Matérn kernel with smoothness 1.5 without nugget. For the bridge and Gaussian process random effect models, we run 11,000 iterations, with the first 1,000 samples discarded as burn-in and 1,000 samples are saved with 10 thin-in rates. The running time of Markov chain Monte Carlo based on a pure R implementation is about 38 mins for the bridge process random effect model and about 20 mins for the Gaussian process random effect model under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
@@
Table 1: Comparison of population-averaged and site-specific estimates across three methods along with coverage probabilities at nominal level 0.95.
Intercept estimate Fixed effect estimate
$\hat{\beta}_{0}^{M}$ $\hat{\beta}_{0}$ $\hat{\beta}_{1}^{M}$ $\hat{\beta}_{1}$
Data generation SpGEE BrP BrP GP SpGEE BrP BrP GP
Bridge process $\rho=0.05$ Bias 0.001 -0.003 -0.021 -0.022 0.015 0.005 0.004 0.002
RMSE 0.237 0.245 0.413 0.412 0.176 0.128 0.068 0.068
Coverage 91.5% 96.0% 96.0% 96.0% 40.0% 96.5% 96.5% 95.5%
Bridge process $\rho=0.1$ Bias -0.013 -0.002 -0.019 -0.019 0.033 -0.009 0.001 -0.001
RMSE 0.386 0.382 0.621 0.617 0.184 0.140 0.074 0.074
Coverage 85.5% 97.5% 97.5% 97.0% 48.7% 95.0% 94.5% 93.5%
Gaussian copula $\rho=0.05$ Bias -0.005 -0.005 -0.007 -0.005 0.008 0.006 0.001 -0.001
RMSE 0.227 0.237 0.339 0.338 0.063 0.051 0.063 0.063
Coverage 92.0% 96.5% 96.5% 95.5% 94.5% 100.0% 98.5% 98.5%
Gaussian copula $\rho=0.1$ Bias -0.011 0.003 0.008 0.006 0.026 -0.021 0.001 -0.002
RMSE 0.380 0.377 0.568 0.568 0.084 0.069 0.065 0.065
Coverage 85.0% 99.0% 99.0% 98.5% 92.0% 100.0% 98.5% 97.0%
@@
SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, bridge process random effect model; GP, Gaussian process random effect model; RMSE, root mean squared error.
@@
The simulation results are summarized in Table 1 in terms of population-averaged estimate $\hat{\beta}^{M}=(\hat{\beta}^{M}_{0},\hat{\beta}^{M}_{1})^{\mathrm{\scriptscriptstyle T}}$ and site-specific estimate $\hat{\beta}=(\hat{\beta}_{0},\hat{\beta}_{1})^{\mathrm{\scriptscriptstyle T}}$, where posterior mean estimates are used for Bayesian approaches. All methods yield consistent estimates in general, but exhibit some differences in terms of mean square error and coverage probabilities depending on underlying data-generating models.
@@
Under the bridge process data-generating model, the spatial generalized estimating equation method gives generally higher mean squared error and narrower confidence intervals of $\hat{\beta}^{M}$, leading to low coverage probability, especially under 50% for $\hat{\beta}^{M}_{1}$. The intercept coverage probability of the spatial generalized estimating equation is slightly lower than the nominal level as the spatial dependence increases, which is consistent with the findings of Cattelan and Varin, (2018). Here the robust standard error correction cannot be easily applied due to the lack of independent replicates; see Cattelan and Varin, (2018) for a discussion. The proposed bridge process random effect model performs well for both $\hat{\beta}^{M}$ and $\hat{\beta}$ as the model is correctly specified, and the Gaussian process random effect model gives a similar site-specific estimate to the proposed model.
@@
Under the Gaussian copula process data-generating model, the spatial generalized estimating equation method provides reasonable coverage probabilities compared to the proposed model, yet has a slightly higher mean square error. The proposed model gives a wider credible interval for $\hat{\beta}^{M}_{1}$, and a similar pattern can be observed for $\hat{\beta}_{1}$ but to a milder degree. In all settings, the mean square error of estimators increases as spatial dependence becomes stronger.
Table 2: Comparison of predictive performance (both higher the better) across three methods based on held-out binary data, averaged across 200 replicates.
Data generation SpGEE BrP GP
Bridge process $\rho=0.05$ Test log-likelihood ×$\times 10^{2}$63.446 -53.073 -53.069
Test AUC ×$\times 10^{2}$ 68.304 80.130 80.137
Bridge process $\rho=0.1$ Test log-likelihood ×$\times 10^{2}$62.390 -49.963 -49.963
Test AUC ×$\times 10^{2}$ 68.541 81.931 81.932
Gaussian copula $\rho=0.05$ Test log-likelihood ×$\times 10^{2}$63.545 -52.564 -52.572
Test AUC ×$\times 10^{2}$ 68.280 80.866 80.861
Gaussian copula $\rho=0.1$ Test log-likelihood ×$\times 10^{2}$62.561 -49.108 -49.103
Test AUC ×$\times 10^{2}$ 68.520 83.033 83.037
Test AUC, area under the receiver operating characteristic curve based on test dataset; SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, bridge process random effect model; GP, Gaussian process random effect model. The Monte Carlo standard errors are all less than 0.5 for $\rho=0.05$ and all less than 0.7 for $\rho=0.1$.
Table 2 summarizes predictive performance based on held-out data of size $n_{\mathrm{test}}=50$, $N_{\mathrm{test}}=500$. We compare test log-likelihood averaged over $N_{\mathrm{test}}=500$ binary data and area under the receiver operating characteristic curve based on test data (test AUC). For the spatial generalized estimating equation method, we used $\mathrm{logit}^{-1}(x_{ij}^{{\mathrm{\scriptscriptstyle T}}}\hat{\beta}^{M})$ as predicted probabilities and for Bayesian approaches, we used the posterior predictive mean. The results suggest that the bridge process and Gaussian process random effect models produce virtually identical predictions which are superior to marginal methods that cannot yield site-specific predictions.
4.2 Scalability analysis with low-rank dependence structure
Next, we conduct another simulation study to analyze the benefits of scalable computing strategies described in Section 3.2 when the number of spatial locations $n$ becomes large. In the random effects model of the bridge process, we compare how the choice of the correlation kernel $\mathcal{R}$ affects the inference, prediction and sampling efficiency of the parameters.
We generate data based on the bridge process random effects model (4). We choose spatial locations uniformly at random from the square domain $\mathscr{S}=[0,2]^{2}$ to decide training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})\in\{(200,50),(800,200)\}$. Similarly to previous simulation settings, we set $\phi=0.7$ and use the Matérn correlation kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where the true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. We set $N_{i}=10$ for all locations, set $p=2$ including the intercept, with $x_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,1)$. The true fixed effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$, and the corresponding true population-averaged coefficient is $\beta^{M}=(\beta_{0}^{M},\beta_{1}^{M})^{\mathrm{\scriptscriptstyle T}}=(0,0.7)^{\mathrm{\scriptscriptstyle T}}$. Denoting $N=\sum_{i=1}^{n}N_{i}$, we have training data sizes $(n,N)\in\{(200,2000),(800,8000)\}$ with two different ranges $\rho\in\{0.05,0.1\}$, and this data generation process is repeated 200 times.
We consider 3 different correlation kernels, one with a full-rank Matern kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ and two with a low-rank kernel (8) with $q=49$ and $q=100$. Within the domain $[0,2]^{2}$, the knot locations $\{\tilde{s}_{k}\}_{k=1}^{q}$ are selected as $\{0.1,0.4,\dots,1.6,1.9\}^{2}$ for $q=49$ and $\{0.1,0.3,\dots,1.7,1.9\}^{2}$ for $q=100$, respectively. The range parameter is assumed to be unknown and we adopt the same prior specification as in Section 4.1. For the Markov chain Monte Carlo, we run 6,000 iterations, with the first 1,000 samples discarded as burn-in and 5,000 samples are saved without thinning.
Figure 2: Boxplot summaries of scalability analysis simulation result for $\rho=0.05$ (moderate spatial dependence) based on 200 replicated datasets.
Figure 3: Boxplot summaries of scalability analysis simulation result for $\rho=0.1$ (strong spatial dependence) based on 200 replicated datasets.
The simulation results are summarized in Figure 2 and Figure 3. In terms of predictive performance based on average test log-likelihood and test AUC, the full-rank correlation kernel performs better than the low-rank ones for $\rho=0.05$ but similarly for $\rho=0.1$ and $q=100$. This is an expected result, since low-rank methods typically yield predictions that smooth over small-scale patterns (Datta et al.,, 2016), thus the difference is more emphasized when $\rho=0.05$. We also compare the sampling efficiency of $\beta^{M}$ and $\beta$ in terms of multivariate effective sample size (Vats et al.,, 2019) divided by running time (mESS/sec). As the number of spatial locations increases from $n=200$ to $n=800$, the mESS/sec between full-rank and low-rank differs in an order of magnitude for both $\beta^{M}$ and $\beta$, clearly showing the computational benefits of the low-rank correlation kernel. The difference is mainly attributed to the running time difference, as mESS values not standardized by time are similar.
5 Childhood malaria prevalence in Gambia